HUBBLE SPACE TELESCOPE DAILY REPORT #2625 PERIOD COVERED: 0000Z (UTC) 05/16/00 - 0000Z (UTC) 05/17/00 Daily Status Report as of 138/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.2 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.3 Completed STIS/CCD 8236 (Black Holes and Gas Disks in a Complete Sample of Radio-Loud Ellipticals - II: Kinematics) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the radio-loud elliptical galaxy UGC-7115. From these observations we will measure the black hole masses, determine the nature, kinematics, and structure of the central gas disks and obtain a modeled upper limit to their mass accretion rates. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.4 Completed Three Sets of STIS/CCD 8808 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted, parallel observations as part of a POMS test proposal. The observations completed with no anomalous activity. 1.5 Completed STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.6 Completed WF/PC-2 8609 (Host Galaxy Luminosities of the Most Luminous QSOs) The WF/PC-2 was used to determine the factors that decide whether a galaxy is host to a luminous QSO. Recent evidence indicates that host galaxy spheroid mass is a key factor: we want to test the relationship between host galaxy and QSO luminosity at the highest feasible QSO luminosities by measuring the luminosities and profiles of host galaxies for QSOs with -25.8 > M_V > -27.6. Current data in this area are conflicting. The observations completed with no reported anomalies. 1.7 Completed STIS/CCD 8489 (Searching for the Hydrogen Reionization Edge of the Universe at 5HST) The Space Telescope Imaging Spectrograph (CCD) was used to observe 36 parallel orbits {4--5 fields of 5--8 orbits each} to constrain the H Lyman-edge in emission that marks the transition from a neutral to a fully ionized IGM. This edge is due to recombination from the H Lyman series and Lyman continuum, and can be used to constrain zion, one of the most important unknown quantities in large scale structure and cosmology. The proposal completed nominally. 1.8 Completed FGS 8292 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) The Fine Guidance Sensor (FGS1R) was used to make astrometric observations of the star GJ469 and other nearby reference stars. This is part of a study to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 8166 (Narrow-Band HAlpha Imaging of the Edge-on Galaxy NGC4631) The WF/PC-2 was used to make observations of nearby edge-on disk galaxies which are essential to determine the role that the disk/halo interaction has in galaxy evolution. Located in a direction of exceptionally low Galactic extinction, NGC4631 is particularly suitable for a multi-wavelength investigation of extra-planar gas components. The observations completed nominally. 1.10 Completed STIS/CCD 8549 (POMS Test Proposal: Non-Scripted Parallel Proposal Continuation I) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.11 Completed STIS/CCD/MA2 8197 (Detection of ^11B/^10B: Part II) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to make spectrographic observations of star HD940287 to measure the boron isotope ratio (i.e., 11B/10B). A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.12 Completed Two Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working. During the second iteration of this proposal, the loss of lock described in 2.1 and HSTAR 7605 occurred. Otherwise, the proposal completed with no further reported problems. 1.13 Completed WF/PC-2 8459 (Decontaminations and Associated Observations Pt. 2/2) The WF/PC-2 was used to perform the monthly decons. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, and UV throughput checks. The observations were completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 16 Successful: 16 Per HSTAR 7605, there was a loss of lock involving FGS-3 at 137/235720Z, after a successful acquisition. Lock was recovered by the issuance of the new #44 commands. The proposal described in 1.12 was affected, as was WF/PC-2 8460 which occurs during the period of the next report. 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared once. The STIS MCE-2 reset at 137/201954Z while the high voltage was off and while the vehicle was outside any SAA interval. The STIS flight software error counter was reset at 137/2158Z as directed by ROP NS-12. MAMA-2 was recovered at 137/2347Z via normal SMS commanding at the time of the following high voltage on event. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW